I have exams this week and there are some things I am not sure what to do for my algebra class I have never been good at probability for example
Eight cards are marked from 1 through 8. What is the Probability of drawing 2 or an odd number?

Answer by jim_thompson5910(35256)

P(drawing 2) = 1/8

P(drawing an odd number) = (# of odd numbers)/(# of total numbers) = 4/8 = 1/2

------------------------

P(drawing 2 or drawing odd number) = P(drawing 2) + P(drawing an odd number)

P(drawing 2 or drawing odd number) = 1/8 + 1/2

P(drawing 2 or drawing odd number) = 1/8 + 4/8

P(drawing 2 or drawing odd number) = 5/8


So the answer is 5/8

If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim